OpenStudy (anonymous):

what is the volume of the region bounded by y=e^-x^2, y=0, x=0, and x=2 when it is rotated around the y axis?

OpenStudy (turingtest):

\[y=e^{-x^2}\]?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

I don't really see how to integrate this is this how the problem was given originally? or did you get this function from somewhere?

OpenStudy (anonymous):

ya thats the equations we were given :/

OpenStudy (dumbcow):

 the volume is sum of the cross-sectional areas from 0 to 1. each area is a circle with radius of x. From y=0 to e^-4, the radius is constant at 2 y = e^-x^2 lny = -x^2 -ln(y) = x^2 Area = pi*r^2 = pi*x^2 = -pi*ln(y) \[\large V = \pi \int\limits_{0}^{e^{-4}} 4 dy + \pi \int\limits_{e^{-4}}^{1}-\ln (y) dy\]